OpenStudy (anonymous):

A cube has all of its sides painted in red and then is cut into 216 cubes of the same size. All cubes are placed in an urn and are thoroughly mixed, so that the probability of being randomly picked from the urn is the same for all cubes. What is the probability that a randomly picked cube has at least one of its sides painted red?

OpenStudy (cwrw238):

the cube has dimensions 6*6*6 so you can calculate how many small cubes are on the outside of the large cube. This number / 216 will give you the required probability

OpenStudy (mathmale):

Hint: any two parallel faces of the original cube have a total of 72 cubes painted red. How many cubes between these two parallel faces also have at least one side painted red?

OpenStudy (cwrw238):

there are 36 cubes on top and bottom = 72 and the rest are in the 4 sides 24 +24 + 16 + 16 = 48 +32 = 80 total 'outside' cubes = 152

OpenStudy (cwrw238):

I think!!!!

OpenStudy (cwrw238):